# Sample Compression Unleashed: New Generalization Bounds for Real Valued Losses

### v2

@@ C.2 Corollaries to the main result
To prove most corollaries, we are going to need the following lemma :
@@
To prove this corollary, we need to do two things : bound $\operatorname{\mathcal{E}}_{\Delta_{C}}$ and rearrange the terms. Both of these things were already proven multiple time in the PAC-Bayes theory, so we simply restate their proofs.
We start by bounding $\operatorname{\mathcal{E}}_{\Delta_{C}}$. We follow the proof of Germain et al., (2015). Let us introduce a random variable $X_{\operatorname{\mathbf{i}}}^{\sigma}$ that follows a binomial distribution of $m$ trials with a probability of success $\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{i}}},\sigma))$, denoted $B(m,\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{i}}},\sigma)))$. We use Lemma 12 with $g(\cdot)=e^{m\Delta_{C}(\cdot,\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(h))}$.
@@
where the last lign is derived using binomial theorem.
@@
To prove this corollary, we need to bound $\operatorname{\mathcal{E}}_{\mathrm{kl}}$. This was first proven by Langford and Seeger, (2001); Seeger, (2002) and then improved by Maurer, (2004). We restate the proof for completeness.
@@
The last two inequalities were proven by Maurer, (2004). The last inequality holds only for $n-\operatorname{|\mathbf{i}|}\geq 8$, but it can be verified that $\operatorname{\mathcal{E}}_{\mathrm{kl}}(\operatorname{\mathbf{i}},\sigma)\leq 2\sqrt{n-\operatorname{|\mathbf{i}|}}$ holds for $n-\operatorname{|\mathbf{i}|}\geq 1$. ∎
@@
We assume that the loss $\ell$ is $\varsigma^{2}$-sub-Gaussian, which is defined as:
@@
In Equation 10, we use the i.i.d. assumption. In Equation 11, we use the $\varsigma^{2}$-sub-Gaussian assumption.
@@ C.3 Behavior with zero error
See 8
Before proving the result, we remind the reader of the definition of the binomial tail:
$\textrm{Bin}(k,m,r)=\sum_{i=0}^{k}\matrixquantity(m\\ i)r^{i}(1-r)^{m-i}$
and the binomial tail inversion :
$\overline{\textrm{Bin}}(k,m,\delta)=\operatorname*{arg\,sup}_{r\in[0,1]}\left\{\textrm{Bin}(k,m,r)\geq\delta\right\}$
Proof of Eq. 3.
@@
Thus, we have :
@@
Proof of Eq. 4.
We want to show that :
$1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})=\lim_{C\to\infty}\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})]$
We simply take the limit.
@@
Proof of Eq. 5.
We want to prove that :
$\lim_{C\to\infty}\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})]=\inf_{C>0}\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})].$
First of all, we know that :
$\inf_{C>0}\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})]=\quantity[1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})]\inf_{C>0}\frac{1}{1-e^{-C}}$
We show that $f(C)=\frac{1}{1-e^{-C}}$ is decreasing on $[0,\infty)$ and has a minimum at $C=\infty$.
We take its derivative :
@@
$=(-1)\frac{e^{-C}}{\quantity(1-e^{-C})^{2}}$
The derivative of $f$ is always negative, so $f$ is decreasing. Thus, it will have no minimum on $\operatorname{\mathbb{R}}_{>0}$. Indeed, if we try to find a value of $C\in\operatorname{\mathbb{R}}_{>0}$ such that $f^{\prime}(C)=0$, we have
$(-1)\frac{e^{-C}}{\quantity(1-e^{-C})^{2}}$ $=0$
@@
There is no $C\in\operatorname{\mathbb{R}}_{>0}$ such that $e^{-C}=0$, however we know that $\lim_{C\to\infty}e^{-C}=0$. This finishes the proof. ∎
Proof of Eq. 6.
We want to show :
$1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})=\mathrm{kl}^{-1}\quantity(0,\frac{1}{m}\ln\frac{1}{\delta})$
The function $\mathrm{kl}(0,p)$ is monotonically increasing and $\mathrm{kl}(0,1)=\infty$. Thus, there exists a value $p^{*}=\mathrm{kl}^{-1}\quantity(0,\frac{1}{m}\ln\frac{1}{\delta})$ such that $\mathrm{kl}(0,p^{*})=\frac{1}{m}\ln\frac{1}{\delta}$.
$p^{*}=\mathrm{kl}^{-1}\quantity(0,\frac{1}{m}\ln\frac{1}{\delta})$
$\iff$ $\mathrm{kl}(0,p^{*})=\frac{1}{m}\ln\frac{1}{\delta}$
$\iff$ $0\ln\frac{0}{p^{*}}+(1)\ln\frac{1}{1-p^{*}}=\frac{1}{m}\ln\frac{1}{\delta}$ (12)
$\iff$ $\ln\frac{1}{1-p^{*}}=\frac{1}{m}\ln\frac{1}{\delta}$
$\iff$ $\frac{1}{1-p^{*}}=\exp\quantity(\frac{1}{m}\ln\frac{1}{\delta})$
$\iff$ $p^{*}=1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})$
In Eq. 12, we use the convention of the PAC-Bayes theory that $00\ln 0=$ 0. This is used to define the $\mathrm{kl}$ when $q\in\{0,1\}$ or $p\in\{0,1\}$. ∎
Proof of Eq. 7.
We want to prove that
$\mathrm{kl}^{-1}\quantity(0,\frac{1}{m}\ln\frac{1}{\delta})\leq\mathrm{kl}^{-1}\quantity(0,\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta}).$
@@
then $p^{*}\leq p^{\dagger}$. ∎
Proof of Eq. 8.
We want to prove that
$\mathrm{kl}^{-1}\quantity(0,\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta})=\mathrm{kl}^{-1}\quantity(0,\frac{1}{m}\ln\frac{1}{\delta})+K(m,\delta)$
and that $K(m,\delta)$ decreases for $m$ large enough and tends to 00 when $m\to\infty$. To do so, we will use Eq. 6, restated here for ease of reading
$1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})=\mathrm{kl}^{-1}\quantity(0,\frac{1}{m}\ln\frac{1}{\delta}).$
We start by defining the constant as the gap between the two terms.
@@
We now prove that this constant is decreasing for $m$ large enough. To do so, we start by computing the derivative of $K$, for $m>0$.
We have :
$\partialderivative{K(m,\delta)}{m}$ $=\frac{1}{2}\frac{\exp\quantity(-\frac{1}{m}\ln\quantity(\frac{2\sqrt{m}}{\delta}))-2\ln(\delta)\exp\quantity(\frac{1}{m}\ln(\delta))-2\ln\quantity(\frac{2\sqrt{m}}{\delta})\exp\quantity(-\frac{1}{m}\ln\quantity(\frac{2\sqrt{m}}{\delta}))}{m^{2}}$
$=\frac{1}{2}\frac{\quantity(1-2\ln\quantity(\frac{2\sqrt{m}}{\delta}))\exp\quantity(-\frac{1}{m}\ln\quantity(\frac{2\sqrt{m}}{\delta}))+2\ln(\frac{1}{\delta})\exp\quantity(-\frac{1}{m}\ln(\frac{1}{\delta}))}{m^{2}}$
The first term is always negative, as we have
$1-2\ln\quantity(\frac{2\sqrt{m}}{\delta})\leq 0\iff\frac{\delta^{2}}{4}e\leq m$
and $\frac{\delta^{2}}{4}e$ is always smaller than one. All the other terms will always be positive.
We focus only on the numerator, as the denominator will always be greater than 00. We wish to find the values of $m$ such that the derivative of $K(m,\delta)$ is negative, showing that $K(m,\delta)$ is decreasing.
$\quantity(1-2\ln\quantity(\frac{2\sqrt{m}}{\delta}))\exp\quantity(-\frac{1}{m}\ln\quantity(\frac{2\sqrt{m}}{\delta}))+2\ln(\frac{1}{\delta})\exp\quantity(-\frac{1}{m}\ln(\frac{1}{\delta}))\leq 0$
$\iff$ $\quantity(1-2\ln\quantity(\frac{2\sqrt{m}}{\delta}))\exp\quantity(-\frac{1}{m}\ln\quantity(2\sqrt{m}))\exp\quantity(-\frac{1}{m}\ln\quantity(\frac{1}{\delta}))+2\ln(\frac{1}{\delta})\exp\quantity(-\frac{1}{m}\ln(\frac{1}{\delta}))\leq 0$
$\iff$ $\quantity(1-2\ln\quantity(\frac{2\sqrt{m}}{\delta}))\exp\quantity(-\frac{1}{m}\ln\quantity(2\sqrt{m}))+2\ln(\frac{1}{\delta})\leq 0$
$\iff$ $2\ln(\frac{1}{\delta})\leq\quantity(2\ln\quantity(\frac{2\sqrt{m}}{\delta})-1)\exp\quantity(-\frac{1}{m}\ln\quantity(2\sqrt{m}))$
$\iff$ $2\ln(\frac{1}{\delta})\leq\quantity(2\ln\quantity(\frac{1}{\delta})+\ln\quantity(2\sqrt{m})-1)\exp\quantity(-\frac{1}{m}\ln\quantity(2\sqrt{m}))$
As the exponential term tends to one when $m$ tends to $\infty$, for $m$ large enough, the inequality will always hold and $K(m,\delta)$ will be decreasing. To the best of our knowledge, there exists no analytical solution to this inequality. However, we can easily compute $m^{*}$ the solution of $\partialderivative{K(m,\delta)}{m}=0$. Then, we know that for any $m\geq m^{*}$, $K(m,\delta)$ will be decreasing. [^2] In Fig. 3, we present the value of $m^{*}$ for multiple values of $\delta$. In Fig. 4, we present the maximum value of $K(m,\delta)$, which is achieved at $m^{*}$, for multiple values of $\delta$.
Figure 3: Illustration of the behavior of the solution $m^{*}$ for different values of $\delta$.
Figure 4: Illustration of the behavior of $K(m^{*},\delta)$ for different values of $\delta$.
Figure 5: Illustration of the behavior of $K(m,\delta)$ for different sizes of datasets, when $\delta=0.01$. We highlight the values corresponding to the datasets used for classification.
Finally, we show that this constant $K(m,\delta)$ tends to 00 when $m$ tends to $\infty$, as such
$\lim_{n\to\infty}K(n,\delta)$ $=\lim_{n\to\infty}\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})-\exp\quantity(-\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta})$
$=\lim_{n\to\infty}\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})-\lim_{n\to\infty}\exp\quantity(-\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta})$
@@
To provide an intuition of the magnitude of $K(m,\delta)$, hence of the tightness of the upper bound of the binomial tail inversion, in Fig. 5 we report the values of $K(m,\delta)$ for different sizes of datasets $m$ and the commonly chosen $\delta=0.01$. In particular, we highlight the values corresponding to the training set sizes used in our experiments, as reported in Table 6. Note however, that when computing the bounds in practice, the gap will be $K(m-\operatorname{|\mathbf{i}|},\delta)$ as we will need to consider the size of the compression set.

### v3

@@ C.2 Corollaries to the main result
To prove most corollaries, we are going to need the following lemma.
@@
The proof is divided in two steps: bounding $\operatorname{\mathcal{E}}_{\Delta_{C}}$ and rearranging the terms. Note that these both steps are common in the proofs of PAC-Bayesian literature. Our proof mainly follow the one of Germain et al., (2015).
We start by bounding $\operatorname{\mathcal{E}}_{\Delta_{C}}$. Let us introduce a random variable $X_{\operatorname{\mathbf{i}}}^{\sigma}$ that follows a binomial distribution of $m$ trials with a probability of success $\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{i}}},\sigma))$, denoted $B(m,\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{i}}},\sigma)))$. We use Lemma 13 with $g(\cdot)=e^{m\Delta_{C}(\cdot,\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(h))}$.
@@
where the last line is derived using binomial theorem.
@@
To prove this corollary, we need to bound $\operatorname{\mathcal{E}}_{\mathrm{kl}}$. In the PAC-Bayes literature, this was first done by Langford and Seeger, (2001); Seeger, (2002) and then improved by Maurer, (2004). We restate the proof of the latter for completeness.
@@
The last two inequalities were proven by Maurer, (2004) for $n-\operatorname{|\mathbf{i}|}\geq 8$. As noticed afterward by Germain et al., (2015), it can be verified numerically that $\operatorname{\mathcal{E}}_{\mathrm{kl}}(\operatorname{\mathbf{i}},\sigma)\leq 2\sqrt{n-\operatorname{|\mathbf{i}|}}$ also holds for $81\leq n-\operatorname{|\mathbf{i}|}<$ 8 .. ∎
@@
We assume that the loss $\ell$ is $\varsigma^{2}$-sub-Gaussian, which is defined as
@@
Equation (9) relies on the i.i.d. assumption and Equation (10), on the $\varsigma^{2}$-sub-Gaussian assumption.
@@ C.3 Behavior with zero error
Let us recall the definition of the binomial tail,
$\textrm{Bin}(k,m,r)=\sum_{i=0}^{k}\matrixquantity(m\\ i)r^{i}(1-r)^{m-i},$
and the binomial tail inversion,
$\overline{\textrm{Bin}}(k,m,\delta)=\operatorname*{arg\,sup}_{r\in[0,1]}\left\{\textrm{Bin}(k,m,r)\geq\delta\right\}.$
The proofs of Theorem 8 and Theorem 9 make use of the following lemma.
Lemma 14.
In the consistent case, the binomial tail has the following analytical solution:
$\overline{\mathrm{Bin}}(0,m,\delta)=1-\exp\quantity(-\frac{1}{m}\ln\quantity(\frac{1}{\delta})).$
Proof.
We start by rewriting the binomial tail distribution, with $k=0$.
@@
We now rewrite the binomial tail inversion.
@@
At Eq. 11, we use the fact that the maximum value of $r$ such that $1-\delta^{\frac{1}{m}}\geq r$ is simply $1-\delta^{\frac{1}{m}}=r$. ∎
We did not find any mention of Lemma 14’s equality in the literature. However, the following inequality is well known and can be found in papers such as Laviolette et al., (2005):
$\overline{\textrm{Bin}}(k,m,\delta)\leq 1-\exp\quantity(\frac{-1}{m-k}\quantity[\ln\matrixquantity(m\\ k)+\ln\quantity(\frac{1}{\delta})]).$
When $k=0$, it reduces to
$\overline{\textrm{Bin}}(0,m,\delta)\leq 1-\exp\quantity(\frac{-1}{m}\ln\quantity(\frac{1}{\delta})).$
We now prove Theorem 8. See 8
Proof.
From Lemma 14, we have
$\overline{\textrm{Bin}}(0,m,\delta)=1-\exp\quantity(\frac{-1}{m}\ln\quantity(\frac{1}{\delta})).$
Thus, we will prove the two following equations :
$1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})=\lim_{C\to\infty}\left\{\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})]\right\}$ (12)
and
$\lim_{C\to\infty}\left\{\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})]\right\}=\inf_{C>0}\left\{\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})]\right\}$ (13)
To prove Eq. 12, we simply compute
@@
Let’s now prove Eq. 13. We know that
$\inf_{C>0}\left\{\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})]\right\}=\quantity[1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})]\inf_{C>0}\left\{\frac{1}{1-e^{-C}}\right\}$
We obtain that function $f(C)=\frac{1}{1-e^{-C}}$ is decreasing on $[0,\infty)$, and thus has a minimum at $C\to\infty$, by the fact that its derivative is negative for all $C\geq 0$:
@@
$=(-1)\frac{e^{-C}}{\quantity(1-e^{-C})^{2}}.$
Note that if one tries to find a value of $C\in\operatorname{\mathbb{R}}_{>0}$ such that the derivative $f^{\prime}(C)$ of the above proof is zero, they obtain
$f^{\prime}(C)=(-1)\frac{e^{-C}}{\quantity(1-e^{-C})^{2}}$ $=0$
@@
There is no $C\in\operatorname{\mathbb{R}}_{>0}$ such that $e^{-C}=0$, however we know that $\lim_{C\to\infty}e^{-C}=0$.
Thus, as the function is monotonically decreasing when $C$ increases, given an arbitrarily small $\epsilon>0$, there is always a $C(\epsilon)$ large enough such that
$\frac{1}{1-e^{-C(\epsilon)}}-1\leq\epsilon.$
For example, with $\epsilon=0.01$ and $C(\epsilon)=4.616$, we have $\frac{1}{1-e^{-C(\epsilon)}}-1=0.00999\leq 0.01$. With $\epsilon=10^{-5}$ and $C(\epsilon)=11.513$, we have $\frac{1}{1-e^{-C(\epsilon)}}-1=0.999\times 10^{-5}\leq 10^{-5}$. Thus, it is possible to be arbitrarily tight to $1$, for any $\epsilon>0$.
We now prove Theorem 9. See 9
Proof.
To prove this result, we need to prove the following sequence of equations:
$\overline{\textrm{Bin}}(0,m,\delta)$ $=\mathrm{kl}^{-1}\quantity(0,\frac{1}{m}\ln\frac{1}{\delta})$ (14)
$\leq\mathrm{kl}^{-1}\quantity(0,\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta})$ (15)
We first prove Eq. 14. We already know that
$\overline{\textrm{Bin}}(0,m,\delta)=\inf_{C>0}\left\{\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})]\right\}.$
Moreover, from the proof of Theorem 3 of Letarte et al., (2019), for any constant $A>0$, we know that
$\inf_{C>0}\left\{\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-C\quantity[\frac{k}{m}]-\frac{1}{m}\ln\frac{A}{\delta})]\right\}=\mathrm{kl}^{-1}\quantity(\frac{k}{m},\frac{1}{m}\ln\frac{A}{\delta}).$ (17)
Thus, with $A=1$, we have
$\overline{\textrm{Bin}}(0,m,\delta)=\mathrm{kl}^{-1}\quantity(0,\frac{1}{m}\ln\frac{1}{\delta}).$
We now prove Eq. 15.
@@
then $p^{*}\leq p^{\dagger}$.
Finally, we prove Eq. 16 and show that $K(m,\delta)$ tends to 00 when $m$ tends to $\infty$ and is bounded by
$0\leq K(m,\delta)\leq\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta}.$
We start by defining the constant $K(m,\delta)$ as the gap between the two following terms:
@@
In the second line, we use Eq. 17 with both $A=2\sqrt{m}$ and $A=1$. ∎
We now highlight some properties of $K(m,\delta)$. First of all, we show that $K(m,\delta)\geq 0$ for $m\geq\tfrac{1}{4}$.
$\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})-\exp\quantity(-\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta})\geq 0$
$\iff$ $\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})\geq\exp\quantity(-\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta})$
$\iff$ $-\frac{1}{m}\ln\frac{1}{\delta}\geq-\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta}$ (18)
$\iff$ $\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta}\geq\frac{1}{m}\ln\frac{1}{\delta}$
$\iff$ $2\sqrt{m}\geq 1$ (19)
$\iff$ $m\geq\frac{1}{4}$
In Eq. 18 and Eq. 19, we use the fact that both the exponentials and logarithms are increasing functions. From this, we also know that $K(\tfrac{1}{4},\delta)=0.$ As the parameter $m$ is the size of a dataset, we know that we always have $m\geq 1$ and thus $K(m,\delta)\geq 0$.
Secondly, we show that $K(m,\delta)$ tends to 00 when $m$ tends to $\infty$.
$\lim_{m\to\infty}K(m,\delta)$ $=\lim_{m\to\infty}\left\{\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})-\exp\quantity(-\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta})\right\}$
$=\lim_{m\to\infty}\left\{\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})\right\}-\lim_{m\to\infty}\left\{\exp\quantity(-\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta})\right\}$
@@
Next, we compute a simple upper bound of $K(m,\delta)$, that also tends to 00 when $m\to\infty$.
$K(m,\delta)$ $=\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})-\exp\quantity(-\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta})$
$\leq 1-\exp\quantity(-\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta})$
$\leq\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta}.$
The last line uses the inequality $1-e^{-x}\leq x$.